# Dynamic tariffs-based demand response in retail electricity market under uncertainty

### v5

@@ 3.5 MILP formulation
$\psi_{jt\omega},\psi_{jt\omega}^{C_{min}},\psi_{jt\omega}^{C_{max}},\tau_{t\omega}^{+},\tau_{t\omega}^{-},\tau_{t\omega}^{q},\tau_{t\omega}^{y},\tau_{t}^{P}\}$ is the set of the decision variables for the competitive model. In particular, the complementarity conditions from the consumer’s KKT conditions can be linearized as follows:
@@
where $M^{c}$ and $M^{P}$ are large enough constants.
@@
where $M^{S}$ and $M^{T}$ are large enough constants.
@@
where $M^{S}$ and $M^{T}$ are large enough constants.
@@ 4 Numerical results and discussion
In this section, we provide numerical examples to complement the analytical results in the previous sections.
@@ 4.1 Data
We consider a day-ahead electricity market clearing price from the European Energy Exchange (EEX). The day-ahead electricity market clearing prices are available in one-hour intervals and can be found at EEX (2023), along with other market data. The marginal utility is the price-demand curve where $A_{jt\omega}$ and $B_{jt\omega}$ are price-demand curve intercept and slope, respectively. By utilizing day-ahead clearing prices and quantities data from a specific date, we can calibrate the expected values for utility parameters ($A_{jt\omega}$ and $B_{jt\omega}$) for consumer $j$, at time $t$ under scenario $\omega$. Subsequently, introducing perturbations allows for the consideration of varying levels of flexibility among residential consumers (Zugno et al., 2013). Regarding consumers’ flexibility, $\Delta_{j}^{C_{max}}=0$ indicates no flexibility, whereas $\Delta_{j}^{C_{max}}>0$ signifies demand flexibility of varying degrees.
Although theoretically there is no limit to the number of consumers, we simplify our model by limiting it to three. This limitation to three consumers is sufficient to yield quantitative insights into the model’s behavior because i) the findings can be generalized to representative real-world scenarios, as previously exemplified by Niromandfam et al. (2020) and Zugno et al. (2013). ii) we address the preferences of three consumers by adjusting demand parameters and offering flexibility in response to changes in consumption. Technically, this entails analyzing the scenarios for each consumer both before and after the implementation of the demand response program, effectively treating the three consumers as if we were dealing with six separate consumers. That means, the expected values of parameters $A_{j}$, $B_{j}$, and $\Delta_{j}^{C_{max}}$ are assumed to vary. This variation reflects the diverse reactions of consumers to price signals from the retailer. Furthermore, each consumer reflects distinct behavior such as Consumer 1 demonstrates the lowest slope and greatest flexibility, while Consumer 3 shows the least flexibility, and Consumer 2 lies between these two extremes (as detailed in Table 2). This approach not only aids in understanding the impact of these parameters on market outcomes but also enhances the visualization of the impact of each parameter on each consumer’s market outcomes.
[…]
@@ 4.2 Case studies
We consider six cases for both the MPEC and MILP (equilibrium) models in the numerical simulation:
@@
    Benchmark: This case serves as a reference for the increase or decrease of demand parameters ($A_{jt\omega},B_{jt\omega},\Delta_{j}^{C_{max}}$) calibrated using the input data from Table 2.
@@
    Linear: Parameter $A_{j}$ is adjusted by 25% increase compared to its value in the Benchmark case, while keeping other parameters constant.
@@
    Quadratic: Parameter $B_{j}$ is adjusted by 35% increase compared to its value in the Benchmark while keeping other parameters constant.
@@
    Flexibility: Parameter $\Delta_{j}^{C_{max}}$ is adjusted by 82% increase compared to its value in the Benchmark case while keeping other parameters constant.
@@
    No flexibility: Parameter $\Delta_{j}^{C_{max}}$ is set to 00.
@@
    Linear utility: parameter $B_{jt\omega}$ is set to 00 and the other parameters kept constant to the Benchmark values.
@@
Based on these six cases, we illustrate the market outcomes for the proposed models, comparing the results across these case studies.
@@
To accommodate the stochastic nature of input parameters in our model, we implement stochastic programming. This approach involves considering scenarios that are assumed to be equiprobable and applying a Gaussian distribution to each parameter. We then utilize the mean values in Table 2 to calibrate simulation data, ensuring the accuracy and reliability of our results. To this end, we select coefficients of variation (CVs) to determine the corresponding standard deviations. For instance, to compute the standard deviation of spot prices, we use the formula $CV\times E[P_{t\omega}^{S}]$, where $E[P_{t\omega}^{S}]$ represents the expected spot price at time $t$ under scenario $\omega$, and we have set the CV at $0.015$ for spot market calculation. Here, $E[P_{t\omega}^{S}]$ represents the day-ahead clearing prices for 24 hours under scenario $\omega$. Following this, $P_{t\omega}^{S}$—the spot price under scenario $\omega$—is used in the simulation to reflect $\omega$ possible scenario regarding market clearing prices. Similarly, we generate $A_{jt\omega}\sim N(\mu_{A_{jt\omega}},\sigma_{A_{jt\omega}})$ and $B_{jt\omega}\sim N(\mu_{B_{jt\omega}},\sigma_{B_{jt\omega}})$, where their standard deviations are calculated using CV values of 0.013 and 0.0013, respectively. Finally, we allow the demand flexibility parameter $\Delta_{j}^{C_{max}}$ to vary from 0 to 5.0 kWh. This range is chosen as it reflects realistic levels of flexibility for individual consumers participating in DR programs, as supported by our analysis and existing literature.
The models have been solved using JuMP version 1.11.1 (Lubin et al., 2023) under the open-source Julia programming Language version 1.8.5 (Bezanson et al., 2017). We use Artelys KNITRO solver version 13.2 (Artelys KNITRO, 2021) for the MPEC and Gurobi version 1.0.1 (Gurobi Optimization, 2021) for the MILP on a CPU E5-1650v2@3.50GHz and 64.00 GB of RAM running workstation. For the MPEC problem, we test the case study with $\Omega=30$ scenarios, and for the MILP problem $\Omega=300$. However, in the MPEC model, the number of scenarios increases, and the number of variables and constraints in the MPEC problem grows exponentially. This growth can potentially render the problem computationally intractable, especially when the number of scenarios exceeds 30($\Omega>30$) [^2] We do not present computational considerations by varying the number of scenarios for the sake of space.. There are possible alternatives to tackle this computational challenge such as utilizing high-performance computing (HPC) with parallel computing techniques (we refer the interested reader to Ahmadi-Khatir et al. (2013) and Nasri et al. (2015)). In our case, we employed the option: KNITRO multistart in parallel with 12 threads. This strategy discards low-quality solutions and initiates the algorithm at different random locations to select the best solutions, should any other local solutions be found. Moreover, a sensitivity analysis on the number of scenarios shows that increasing the number of scenarios has no significant impact on the reported market outcomes (see Table 3). It should also be noted that in the case of the MILP, the computational issue is considerably less concerning. For instance, the model can be efficiently solved even for larger sets of scenarios, such as $\Omega=500$ or greater. The MPEC takes $42.422$ CPU time with $\omega=30$, while MILP takes $15.019$ CPU time with $\omega=300$. It is also important to note that the methods to adjust big Ms are mostly heuristics and problem-dependent, where the trial-and-error procedure has been used in most research work (Motto et al., 2005; Garcés et al., 2009; Jenabi et al., 2013). However, for the particular problem addressed, the tune-up of these constants is achieved without particular numerical trouble. Nevertheless, caution must be exercised for the big M values related to the dual variables to ensure that their value is sufficiently large to not impose any additional constraints to the problem, but not too large so that the resulting model is not ill-conditioned (Pineda and Morales, 2019).
@@
Table 3: Market outcomes sensitivity on the number of scenarios in the MPEC model under the Benchmark case.
@@ 4.3 Market outcomes: Electricity price tariffs
The retailer’s price tariffs in both the MPEC and the MILP models are depicted along with the expected day-ahead electricity market prices in Figure 3b(a). The price tariffs are compared across the Benchmark, Flexibility, and No flexibility cases in both models. In all cases, the MILP price tariffs demonstrate consistent patterns with the market clearing prices in the day-ahead market. However, with the MPEC model, when the retailer’s market power is considered, price tariffs exceed the expected day-ahead electricity prices even during low-demand periods. This is market inefficiency, primarily due to market distortion caused by the strategy of the retailer.
@@
On the other hand, within the competitive equilibrium model (MILP), price tariffs are generally observed to be lower than those in the MPEC model across most cases and periods. The overall price tariffs $P_{t}$ set by the retailer under the MPEC framework are consistently higher than the spot market prices $P_{t\omega}^{S}$. Furthermore, the retailer reduces its procurement when the spot market prices are relatively high (at $9:00,10:00,17:00-21:00$). On the other hand, it increases its purchases when the spot market prices are lower (at $t=1:00-8:00,10:00-17:00$). These electricity price patterns align with the quantities of power the retailer purchases during peak periods, which is reflected in the optimal outcomes showing lower quantities in the spot market. A more detailed analysis can be found in Figure 4, where there is a comprehensive comparison of the results both before and after the implementation of the demand response program under MPEC and MILP models. Higher quantities are purchased when wholesale market prices are lower. Although lower electricity price tariffs are already in place during the early hours, the introduction of increased demand flexibility results in a further decrease in electricity price tariffs in the MPEC model. However, in the equilibrium model (MILP), the overall electricity price tariff consistently remains below the spot price, except during peak periods.
@@
The rise in consumer flexibility leads to lower retailer price tariffs in the MPEC, especially after 7:00. This indicates that if consumers adjust their consumption, electricity prices can decrease, and conversely, prices can increase with less flexibility, even in the presence of market distortion caused by strategic retailers. More interestingly, the MPEC model also has the peak-shaving/valley-filling effect; that is, the realized price tariffs at on-peak and off-peak periods are very similar in the case of high flexibility. This is a much-desired result, as a flatter price tariffs curve means that the system is more predictable and reliable. In the MILP model, however, high demand elasticity notably reduces electricity prices and increases the quantities purchased in both spot and retail markets to a greater extent than observed in the MPEC model. Thus, without demand flexibility, price tariffs are higher in the Benchmark under the MPEC model than in the MILP. Prices do not significantly change over time in the Benchmark case compared with case Flexibility case.
@@
In Figure 3b(b), the impact of increasing demand parameters on price tariffs is illustrated. While an increase in $A_{jt\omega}$ slightly increases the clearing price, the overall impact of demand parameters and the penalty weight on price tariffs is insignificant. Note that imposing a penalty on the retailer’s objective function does affect the power imbalance in the periods considered rather than its impact on electricity price tariffs. Interestingly, increasing the flexibility parameter $\Delta_{j}^{C_{max}}$ leads to a significant reduction in price tariffs, more so than other model parameters. Upon analyzing the market outcomes of the proposed models, a key insight is that market power hinders the efficiency of demand flexibility, where the price-maker retailer exploits consumers’ flexibility to maximize its expected profits by imposing higher tariffs.
4.4 Market outcomes: Loads and spot market quantities
@@
Figure 4: Comparison of consumption and spot market quantities with No flexibility and Flexibility cases.
@@
Figure 4(a) depicts the aggregate consumer purchases in the retail market both before and after the implementation of demand response, in the MPEC and MILP models. The results indicate that lower spot market prices lead to increased purchases by consumers in the MILP model. The trend is similar without the implementation of the demand response program. With MILP, a higher value of $B$ (Quadratic case) is associated with significant and consistent electricity purchases in the retail market. Note that an increase in the quadratic term of the utility function is analogous to a decrease in electricity prices. Both parameters are inversely related to the quantity of electricity purchased. In contrast, an increase in the parameter in $A$ (Linear case) encourages consumers to buy more electricity during off-peak periods in both MPEC and MILP models, thereby avoiding higher tariff periods. Consequently, lower prices during off-peak periods prompt higher purchases from the spot and the retail markets. As loads shift from expensive to cheaper periods using consumers’ flexibility, new peak and valley periods emerge. This indicates that a peak period is typically followed by a valley period. Note that the detailed results for Linear and Quadratic cases are omitted for brevity.
@@
With different cases, the retailer buys higher quantities in the perfect competition where there is a lower price, particularly with case Quadratic (higher $B$). With the Benchmark case, where there are lower values for $A$ and $\Delta_{j}^{C_{max}}$, the retailer purchases a lower quantity even in the MILP model. The equilibrium model shows the higher the competition, the lower the impact the retailer could have on the market outcomes. Specifically, the price tariffs sent by the retailer are lower in the MILP model as the retailer is not a price-maker. This enables consumers to enjoy greater benefits when purchasing their electricity in a competitive market model.
@@
Figure 4(b) shows the quantities purchased in the spot market with MPEC and MILP under the Benchmark case. The spot market quantities show consistency in both models before and after implementing the DR program. However, with high Flexibility, the quantities in the spot market during off-peak periods are significantly higher. This shows that consumer flexibility allows retailers to shift their purchase plans to off-peak periods to benefit from lower prices in the wholesale market. During peak demand periods, on the other hand, retailers might opt out of buying from the day-ahead market due to the high spot market prices. If retailers have to purchase electricity during peak periods when wholesale prices are high, they will impose higher price tariffs on consumers during these periods. This, in turn, motivates consumers to shift/shave their consumption, which benefits both parties.
@@
Despite the aggregate purchase trend reflecting individual consumer purchases, a comparison of consumers’ purchases in the Benchmark case reveals distinct consumer behaviors based on their utility function parameters. For instance, Consumer $1$ tends to buy more than other consumers when electricity price tariffs are low, typically in the early morning and late evening. This is because Consumer $1$ has the minimum slope ($A_{1tw}=0.0013$) and the highest demand flexibility ($\Delta_{j}^{C_{max}}=2.50$) in the Benchmark case. Individual consumers’ results are omitted for brevity.

### v6

@@ 3.5 Equilibrium problem formulation
$\psi_{jt\omega},\psi_{jt\omega}^{C_{min}},\psi_{jt\omega}^{C_{max}},\tau_{t\omega}^{+},\tau_{t\omega}^{-},\tau_{t\omega}^{q},\tau_{t\omega}^{y},\tau_{t}^{P}\}$ is the set of decision variables for the competitive model. In particular, the complementarity conditions from the consumer’s KKT conditions can be linearized as follows:
@@
where $M^{c}$ and $M^{P}$ are sufficiently large constants.
@@
where $M^{S}$ and $M^{T}$ are sufficiently large constants.
@@
where $M^{S}$ and $M^{T}$ are sufficiently large constants.
@@ 4 Numerical results and discussion
This section presents numerical simulations and discusses the models’ practical implications. It also analyzes the models’ performance under various market conditions and demand response scenarios, with sensitivity analyses to ensure the robustness of the results. The numerical settings for the simulations, including data sources, parameter values, and scenario generation are also discussed. These settings are based on standard practices in the literature and are validated against real-world data from the European Energy Exchange platform. Additionally, we perform sensitivity analyses on some key model parameters to confirm the robustness and reliability of our findings.
@@ 4.1 Data
We consider a day-ahead electricity market clearing price from the European Energy Exchange (EEX). The day-ahead electricity market clearing prices are available in one-hour intervals and can be found at EEX (2023). The marginal utility is the price-demand curve, where $A_{jt\omega}$ and $B_{jt\omega}$ are the price-demand curve intercept and slope, respectively. By utilizing day-ahead clearing price and quantity data from a specific date, we can calibrate the expected values for utility parameters ($A_{jt\omega}$ and $B_{jt\omega}$) for consumer $j$, at time $t$ under scenario $\omega$. Subsequently, introducing perturbations allows the levels of flexibility among residential consumers to vary (Zugno et al., 2013). Regarding consumers’ flexibility, $\Delta_{j}^{C_{max}}=0$ indicates no flexibility, while $\Delta_{j}^{C_{max}}>0$ signifies consumers’ flexibility to varying degrees.
Although theoretically there is no limit to the number of consumers, we simplify our model by limiting it to three consumers. This limitation to three consumers is sufficient to draw insights from the considered model, and i) the findings can be generalized to representative real-world scenarios (Niromandfam et al., 2020; Zugno et al., 2013). ii) We address the preferences of three consumers by adjusting demand parameters and offering flexibility in response to changes in consumption. This entails analyzing the cases of each consumer both before and after the implementation of the demand response program. This means that expected values of parameters $A_{j}$, $B_{j}$, and $\Delta_{j}^{C_{max}}$ are assumed to vary. This reflects the reactions of consumers to the retailer’s price signals while competing with each other to impress the retailer and by maximizing their utility(minimizing their disutility). Consumer $1$ demonstrates the lowest slope and greatest flexibility, while Consumer $3$ shows the least flexibility, and Consumer $2$ lies between these two extremes (as detailed in Table 2). This helps us to understand the impact of the parameters on market outcomes and their dynamics.
Table 2: Mean values of consumer parameters in the six case studies ($j=1,2,3$): $A_{j}$, $B_{j}$, and $\Delta_{j}^{C_{max}}$ are expressed in units of €/kWh, €/kWh2, and kWh, respectively.
@@ 4.2 Case studies
We consider six cases for both the MPEC and MILP (equilibrium) models in the numerical analysis:
@@
    Benchmark: This case serves as a reference for the increase or decrease in demand parameters ($A_{jt\omega},B_{jt\omega},\Delta_{j}^{C_{max}}$) calibrated using the input data from Table 2.
@@
    Linear: Parameter $A_{j}$ is adjusted by a 25% increase compared to its value in the benchmark case while keeping the other parameters constant.
@@
    Quadratic: Parameter $B_{j}$ is adjusted by a 35% increase compared to its value in the benchmark while keeping the other parameters constant.
@@
    Flexibility: The parameter $\Delta_{j}^{C_{max}}$ is adjusted by an 82% increase compared to its value in the benchmark case while keeping the other parameters constant.
@@
    No flexibility: The parameter $\Delta_{j}^{C_{max}}$ is set to 00.
@@
    Linear utility: The parameter $B_{jt\omega}$ is set to 00, and the other parameters are kept constant at the benchmark values.
@@
Based on the cases, we illustrate and compare the market outcomes across these case studies.
@@
To account for uncertainty, we implement stochastic programming with equiprobable scenarios and apply Gaussian distributions to the parameters. We then utilize the mean values in Table 2 to calibrate the data and select the coefficients of variation (CVs) for the parameters to determine the corresponding standard deviations. For instance, to compute the standard deviation of spot prices, we use the formula $CV\times E[P_{t\omega}^{S}]$. Thus, $E[P_{t\omega}^{S}]$ represents the expected spot price at time $t$ under scenario $\omega$, and we set the CV at $0.015$ for the spot market calculation. $E[P_{t\omega}^{S}]$ is the day-ahead clearing prices for $24$ hours under scenario $\omega$. $P_{t\omega}^{S}$—the spot price under scenario $\omega$—is used for the simulation to reflect $\omega$ possible scenarios regarding market clearing price dynamics. Similarly, we generate $A_{jt\omega}\sim N(\mu_{A_{jt\omega}},\sigma_{A_{jt\omega}})$ and $B_{jt\omega}\sim N(\mu_{B_{jt\omega}},\sigma_{B_{jt\omega}})$, where their standard deviations are calculated using CV values of 0.013 and 0.0013, respectively. Finally, the demand flexibility parameter $\Delta_{j}^{C_{max}}$ can vary from 0 to 5.0 kWh. This range is chosen because it reflects realistic levels of flexibility for individual consumers participating in DR programs.
The models were solved using JuMP version 1.11.1 (Lubin et al., 2023) under the open-source Julia programming language version 1.8.5 (Bezanson et al., 2017). We use Artelys KNITRO solver version 13.2 (Byrd et al., 2006) for the MPEC and Gurobi version 10.1 (Gurobi Optimization, 2023) for the MILP on a CPU E5-1650v2@3.50GHz and 64.00 GB of RAM running workstation. For the MPEC problem, we test the case study with $\Omega=30$ scenarios, and for the MILP problem $\Omega=300$. However, in the MPEC model, the number of scenarios increases, and the number of variables and constraints in the MPEC problem increases exponentially. This renders the problem computationally intractable, especially when the number of scenarios exceeds 30($\Omega>30$) [^1] We do not present computational considerations by varying the number of scenarios for the sake of space.. There are possible alternatives to tackle this computational challenge, such as utilizing high-performance computing (HPC) with parallel computing techniques (we refer the interested reader to Ahmadi-Khatir et al. (2013) and Nasri et al. (2015)). Since our problem is nonconvex as it involves solving the MPEC problem without applying any linearization or approximations, we tackled it by using the NLP solver KNITRO. Although the presented solutions are denoted as optimal by the solver, proving that their global optimality is not analytically possible, we addressed this issue numerically by employing a multistart strategy. This approach involves running the solver from multiple initial points, whereby the algorithm selects the best solutions after discarding potential lower-quality optima.
Moreover, a sensitivity analysis on the number of scenarios shows that increasing the number of scenarios has no significant impact on market outcomes (see Table 3). For the equilibrium problem with MILP, there is no computational issue concerning the number of scenarios. The model can be efficiently solved for scenarios, such as $\Omega=500$ or greater. The MPEC requires $42.422$ of CPU time with $\omega=30$, while the MILP requires $15.019$ of CPU time with $\omega=300$. It is also important to note that the methods to adjust big-Ms are mostly heuristic and problem-dependent, where the trial-and-error procedure has been used in most related research (Motto et al., 2005; Garcés et al., 2009; Jenabi et al., 2013). However, for the particular problem addressed, the tuning of these constants is achieved without particular numerical trouble. The primary challenge lies in selecting appropriate Big-M values associated with the dual variables. These values must be large enough to avoid imposing unnecessary constraints on the problem. However, they must not be excessively large, as that would risk creating an ill-conditioned model (Pineda and Morales, 2019).
@@
Table 3: Market outcome sensitivity to the number of scenarios in the MPEC model under the benchmark case.
@@ 4.3 Market outcomes: Electricity price tariffs
The retailer’s price tariffs in both the MPEC and the MILP models are depicted along with the expected day-ahead electricity market prices in Figure 3b(a). The price tariffs are compared across the benchmark, flexibility, and no flexibility  [^2] Note that we use No flexibility to refer to the scenario before DR implementation and flexibility to describe the case after DR implementation. cases in both models. In all cases, the MILP price tariffs demonstrate consistent patterns with the market clearing prices in the day-ahead market. However, with the MPEC model, when the retailer’s market power is considered, price tariffs exceed the expected day-ahead electricity prices even during low-demand periods. This is market inefficiency, primarily due to market distortion caused by the strategy of the retailer.
@@
On the other hand, within the competitive equilibrium model (MILP), price tariffs are generally lower than those in the MPEC model across most cases and periods. The overall price tariffs $P_{t}$ set by the retailer under the MPEC framework are consistently higher than the spot market prices $P_{t\omega}^{S}$. Furthermore, the retailer reduces its procurement when the spot market prices are relatively high (at $9:00,10:00,17:00-21:00$). It increases its purchases when the spot market prices are lower (at $t=$1:00-8:00,10:00-17:00). These electricity price patterns align with the quantities of power the retailer purchases during peak periods, which is reflected in the optimal outcomes showing lower quantities in the spot market. A more detailed analysis is available in Figure 4. It provides a comprehensive comparison of results before and after implementing the demand response program under the MPEC and MILP models. Higher quantities are purchased when wholesale market prices are lower. Although lower electricity price tariffs are already in place during the early hours, demand flexibility further decreases electricity price tariffs in the MPEC model. In the equilibrium model (MILP), the overall electricity price tariffs remain below the spot price outside the peak periods.
@@
Consumer flexibility results in lower retailer price tariffs in the MPEC, especially after 7:00. This means that if consumers adjust their consumption, they can further reduce their electricity prices. Otherwise, prices can increase with less flexibility. The MPEC model also has a peak-shaving and valley-filling effect. In cases of high flexibility, the realized price tariffs during on-peak and off-peak periods become very similar. This is a much-desired result, as a flatter price tariffs curve means that the system is more predictable and reliable. In the MILP model, however, high demand elasticity reduces electricity prices and increases the quantities purchased in both the spot and retail markets more than in the MPEC model. Thus, without demand flexibility, price tariffs are greater in the benchmark under the MPEC model than in the MILP model. Compared with those in the flexibility case, prices do not significantly change over time in the benchmark case.
@@
Figure 3b(b) illustrates the impact of increasing demand parameters on price tariffs. While an increase in $A_{jt\omega}$ slightly increases the clearing price, the overall impact of demand parameters and the penalty weight on price tariffs is insignificant. Note that imposing a penalty on the retailer’s objective function affects the power imbalance in the periods considered rather than its impact on electricity price tariffs. Interestingly, increasing the flexibility parameter $\Delta_{j}^{C_{max}}$ leads to a significant reduction in price tariffs, more than other model parameters. A key takeaway is that market power hinders the efficiency of consumer flexibility, where the price-maker retailer exploits consumer flexibility to maximize its expected profits by signaling higher tariffs.
4.4 Market outcomes: Loads and spot market qunatities
@@
Figure 4: Comparison of consumption and spot market quantities for the no flexibility and flexibility cases.
@@
Figure 4(a) depicts the aggregate consumer purchases in the retail market both before and after the implementation of the demand response in the MPEC and MILP models. The results indicate that lower spot market prices lead to increased purchases by consumers in the MILP model. The trend is similar without the implementation of the demand response program. With MILP, a higher value of $B$ (quadratic case) is associated with significant and consistent electricity purchases in the retail market. Note that an increase in the quadratic term of the utility function is analogous to a decrease in electricity prices. Both parameters are inversely related to the quantity of electricity purchased. In contrast, an increase in the parameter in $A$ (linear case) encourages consumers to buy more electricity during off-peak periods in both the MPEC and MILP models, thereby avoiding higher tariff periods. Consequently, lower prices during off-peak periods prompt higher purchases from the spot and retail markets. As loads shift from expensive to cheaper periods using consumer flexibility, new peak and valley periods emerge. This indicates that a peak period is typically followed by a valley period. Note that the detailed results for the linear and quadratic cases are omitted for brevity.
@@
With different cases, the retailer buys higher quantities in perfect competition where there is a lower price, particularly in the quadratic case (higher $B$ values). With the benchmark case, where there are lower values for $A$ and $\Delta_{j}^{C_{max}}$, the retailer purchases a lower quantity even in the MILP model. The equilibrium model shows that the greater the competition is, the lower the impact the retailer could have on the market outcomes. Specifically, the price tariffs sent by the retailer are lower in the MILP model because the retailer is not a price-maker. This enables consumers to enjoy greater benefits when purchasing their electricity in a competitive market model.
@@
Figure 4(b) shows the quantities purchased in the spot market with MPEC and MILP under the benchmark case. The spot market quantities are consistent in both models before and after the DR program was implemented. However, with high flexibility, the quantities in the spot market during off-peak periods are significantly greater. This shows that consumer flexibility allows retailers to shift their purchases to off-peak periods to benefit from lower prices in the wholesale market. During peak demand periods, retailers might decide not to buy from the day-ahead market due to the high spot market prices. If retailers have to purchase electricity during peak periods when wholesale prices are high, they can send higher price tariffs to consumers during these periods. This further motivates consumers to shift/shave their consumption, which benefits both parties.
@@
A comparison of consumers’ purchased quantities in the benchmark case shows individual consumer behaviors based on their utility function parameters. Consumer $1$ tends to buy more than other consumers when electricity price tariffs are low, such as in the early morning and late evening. This is because Consumer $1$ has the minimum slope ($A_{1tw}=0.0013$) and the highest demand flexibility ($\Delta_{j}^{C_{max}}=2.50$) in the benchmark case. Individual consumers’ results are omitted for brevity.